OpenStudy (anonymous):

Evaluate. 4) 57 + 19y + -76 5) 7x - 18 = -74 explain pls

OpenStudy (anonymous):

For the first one, did they tell you what y was?

OpenStudy (anonymous):

no. you have to solve for y.

OpenStudy (anonymous):

Not possible; it's not an equation. Did you copy it down correctly?

OpenStudy (anonymous):

o woops its 57 + 9y = -76

OpenStudy (anonymous):

57 + 19y = -76

OpenStudy (cinniomon):

Ok i'll give it a shot

OpenStudy (cinniomon):

First subtract 57 from both sides this gets you 19y=-133 Next divide both sides by 19 this gets you y=-7 as your answer

OpenStudy (cinniomon):

Now for the other equation 7x-18=-74 Add 18 to both sides which gets you 7x=-56 Now divide both sides by 7 and you get x=-8 as your answer

OpenStudy (cinniomon):

Is this a good explaination?

OpenStudy (anonymous):

Yes, thanks.

OpenStudy (cinniomon):

You r welcome if you have any other questions ill do my best to answer